OpenStudy (anonymous):

How do you find the minimum of the graph y=x^2-3x+1? (the points of the minimum)

OpenStudy (anonymous):

use the magic formula \(-\frac{b}{2a}\) to get the first coordinate of the vertex second coordinate is the minimum

OpenStudy (anonymous):

in your case \(-\frac{b}{2a}=-\frac{-3}{2\times 1}=\frac{3}{2}\) replace \(x\) by \(\frac{3}{2}\) and see what you get

OpenStudy (anonymous):

great explanation satallite :D

OpenStudy (anonymous):

i like the magic formula, it is easy to prove real easy to use and only contains five symbols: minus sign, fraction bar, the number 2 and the variables \(a\) and \(b\)

OpenStudy (anonymous):

differentiate the equation and equate to zero

OpenStudy (anonymous):

what means "differentiation"?

OpenStudy (anonymous):

I think that is out of scope for the OP @DHASHNI

OpenStudy (anonymous):

\(y=ax^2+bx+c=a(x+\frac{b}{2a})^2+k\) vertex is clearly at \((-\frac{b}{2a},k)\) since the first term is a perfect square and is therefore either at most zero or at least zero, depending on the sign of \(a\)

OpenStudy (anonymous):

OP?

OpenStudy (anonymous):

Oringal Poster

OpenStudy (anonymous):

you down with OP P?

OpenStudy (anonymous):

y'=2x-3 equating it to zero 2x-3=0 2x=3 x=3/2

OpenStudy (anonymous):

http://www.youtube.com/watch?v=mJgFU3U4X_Y

OpenStudy (anonymous):

so the points are (1.5, -1.25)? (:

OpenStudy (anonymous):

I like OPP

OpenStudy (anonymous):

LOL!!!

OpenStudy (anonymous):

Other peoples property?

OpenStudy (anonymous):

"propery" maybe ...

OpenStudy (anonymous):

@Genuine yes you are right

OpenStudy (anonymous):

YAY! Thank you so much for helping me today @satellite73

OpenStudy (anonymous):

yw